OpenStudy (anonymous):

Find the slope,if it exist 9x-4y+17=0

OpenStudy (anonymous):

slope of a line ax+by+c=0 is -a/b

OpenStudy (anonymous):

can you re-write that so I can understand

OpenStudy (anonymous):

m=?

OpenStudy (anonymous):

slope = m = -b/a = -(-4)/9 = 4/9